can you please help me slove this problem?
6 2\3 times 3 5\6

Answer by checkley71(8403)

(6 2/3)(3 5/6) FIRST YOU MUST CHANGE THESE MIXED FRACTIONS TO IMPROPER FRACTIONS THUS:
6 2/3=(3*6+2)/3=20/3
3 5/6=(6*3+5)/6=23/6
NOW YOU CAN MULTIPLY THEM
(20/3)(23/6)=(20*23)/(3*6)=460/18=25 5/9 ANSWER.